OpenStudy (anonymous):

You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($10 then $1).

OpenStudy (anonymous):

@estudier help lol

OpenStudy (anonymous):

OK, this is more complicated. First off, how many possibilities (like 6 for a dice)?

OpenStudy (anonymous):

looks like 9

OpenStudy (anonymous):

That's right. Now chance of a random draw being a 10?

OpenStudy (anonymous):

2/9 ?

OpenStudy (anonymous):

Right again. Now assume you have gotten the 10 and put it in your pocket. Chance of a 1?

OpenStudy (anonymous):

3/9 ?

OpenStudy (anonymous):

It is no longer 9, is it?

OpenStudy (anonymous):

oh no lol

OpenStudy (anonymous):

it's 3/8?

OpenStudy (anonymous):

That's right. Now we want the probability of both events happening one after the other. So we multiply the probabilities together.

OpenStudy (anonymous):

3/9 times 2/9 times 3/8 ?

OpenStudy (anonymous):

Think you are mixed up, there are only two probabilities The ten, 2/9 and The one, 3/8

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

5/72?

OpenStudy (anonymous):

You have added the top and multiplied the bottom....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Um...why?

OpenStudy (anonymous):

i messed up? lol

OpenStudy (anonymous):

2*3/9*8 = 6/72 = 1/12 or 1 in 12

OpenStudy (anonymous):

oh i forgot to simplify lol

OpenStudy (anonymous):

Did they give you that as an option?

OpenStudy (anonymous):

yes they did :)